40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1450     Entry time: Wed Apr 1 16:14:36 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: 3.8kHz peak does not change with SRC offset 
Yoichi, Peter

We suspected that maybe the 3.8kHz peak is the DARM RSE somehow coupled to the CARM.
So we added an offset to the SRC error signal to see if the peak moves by changing the offset.
It didn't (at least by changing the SRC offset by +/-1000).
(I had a nice plot showing this, but dtt corrupted the data when I saved it. So no plot attached.)

I also played with the PRC, DARM offsets which did not have any effect on the peak.
The only thing, I could find so far, having some effect on the peak is the arm power. As the arm power is increased, the peak height goes up and the frequency shifts slightly towards lower frequencies.
ELOG V3.1.3-